# Construction of graph-directed invariant sets of weak contractions on semi-metric spaces

### v1

@@ Abstract.
We present a construction of graph-directed invariant sets of weak contractions ni the sense of Matkowski on semi-metric spaces. We follow the approach by Bessenyei and Pénzes, which applies the Kuratowski noncompactness measure without relying on Blascke’s completeness theorem. We also establish a relationship between this approach and a generalized de Rham’s functional equation indexed by a finite directed graph.
@@ 1. Introduction
We now assume that a finite directed multigraph allowing self-loops and a metric space are given, with a transformation on the metric space assigned to each directed edge. This setting allows us to define a natural analogue of Hutchinson’s self-similar sets, known as graph-directed fractal sets. These are described as vectors of non-empty compact subsets and can be realized as fixed points of certain transformations on product spaces. See (1.1) below for the precise definition of the transformation. The fixed point is called a graph-directed invariant set, which is a notion first introduced by Mauldin and Williams [20] as a generalization of Hutchinson’s self-similar sets. When the underlying graph consists of a single vertex, the framework reduces to the classical IFS. Similar to the self-similar sets, the graph-directed invariant sets and related topics have been studied extensively. For further details, see the book by Mauldin and Urbański [19].
In this paper, we construct an invariant vector of non-empty compact subsets for graph-directed IFSs consisting of weak contractions on semi-metric spaces. Semi-metric is obtained by omitting the triangle inequality from Fréchet’s definition of metric. We work specifically with regular semi-metrics introduced by Bessenyei and Páles [1]. Informally, the local properties of semi-metric spaces are similar to those of metric spaces, on the other hand, a semi-metric can differ significantly in global structure. We assign weak contractions in the sense of Matkowski [18] as the components of the graph-directed IFS. Since weak contractions are less restrictive than standard contractions, more careful analysis is required. Our arguments depend on the fixed point theorem provided in [1, Theorem 1].
Our approach follows the method introduced by Bessenyei and Pénzes [2, 3, 4], which differs from Hutchinson’s approach. The method of Bessenyei and Pénzes avoids the use of Banach’s fixed point theorem which depends on Blaschke’s completeness theorem [5], and instead applies Kuratowski’s measure of noncompactness [16]. This enables us to approximate the unique invariant non-empty compact subset by an explicitly constructible, increasing sequence of finite sets. Furthermore, it clearly demonstrates that the invariant subset depends on the topology, rather than on the metric.
@@ 1.1. Framework and main result
Let $X$ be a non-empty set. We say that a map $d:X^{2}\to[0,\infty)$ is a a semi-metric on $X$ if $d(x,y)=0$ if and only if $x=y$, and $d(x,y)=d(y,x)$ for every $x,y\in X$. Let $B(x,r):=\{y\in X|d(x,y)<r\}$ for $x\in X$ and $r>0$. We say that a subset $U$ of $X$ is an open set if for every $x\in X$, there exists $r>0$ such that $B(x,r)\subset U$. This defines a topology on $X$. We remark that $B(x,r)$ can be a non-open set. A closed set is a complement of an open set. For $H\subset X$, we denote the closure of $H$ by $\overline{H}$. We say that a sequence $(x_{n})_{n}$ in $X$ is convergent if $\lim_{n\to\infty}d(x_{n},x)=0$ for some $x\in X$, and it is a Cauchy sequence if $\lim_{n,m\to\infty}d(x_{n},x_{m})=0$. We say that $(X,d)$ is complete if every Cauchy sequence is convergent.
We assume a notion which plays a role in the triangle inequality. For $u,v\in[0,\infty)$, let
@@
We say that a semi-metric space $(X,d)$ satisfies the regular property for the basic triangle function if $\Phi_{d}$ is continuous at $(0,0)$, and, it satisfies the normal property for the basic triangle function if $\Phi_{d}(u,v)\neq\infty$ for $u,v\in[0,\infty)$. Every metric space is a semi-metric space satisfying the regular and normal properties for the basic triangle function. The regular property was introduced by [1] and it plays a role of the triangle inequality. Informally, we can say that a regular metric satisfies a kind of the triangle inequality in “small" scales. The normal property was introduced by [4] and it is important when we deal with bounded subsets. The regular or normal property is different from the notion of the regular or normal topological space (${\bf T}_{3}$ space or ${\bf T}_{4}$ space).
@@
We adopt Matkowski’s definition of weak contractions [18]. We say that a map $\phi:[0,\infty)\to[0,\infty)$ is a comparison function if $\phi$ is increasing, and $\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$. Let $\phi$ be a comparison function. We say that a map $T:X\to X$ is a $\phi$-contraction if $d(Tx,Ty)\leq\phi(d(x,y))$ for every $x,y\in X$.
We state the framework of the graph-directed IFSs by [10, Section 3.1]. Let $q\geq 1$ and $V:=\{1,\cdots,q\}$. For $i,j\in V$, let $E_{i,j}$ be a finite set of directed edges from $i$ to $j$. We remark that $i=j$ is allowed. We allow multiple edges and self-loops in order to make the setting here a generalization of the self-similar set, which corresponds to the case that $q=1$. We assume that for every $i\in V$, there exists $j\in V$ such that $E_{i,j}\neq\emptyset$. This assumption is weaker than the transitivity condition. Assume that for each edge $e\in E_{i,j}$, a right-continuous comparison function $\phi_{e}$ and a $\phi_{e}$-contraction $f_{e}:X\to X$ are given. The case that $q=1$ gives the framework of Hutchinson’s self-similar sets.
Let $\mathcal{P}(X)$ be the set of non-empty subsets of $X$. Hereafter, we need to deal with product spaces $X^{q}$ and $\mathcal{P}(X)^{q}$. Elements in those product spaces are written in bold face.
@@
For a family $({\bf H}_{n})_{n}$ in $\mathcal{P}(X)^{q}$, let $\cup_{n}{\bf H}_{n}:=\left(\cup_{n}H_{n,1}),\dots,\cup_{n}H_{n,q}\right)$, where ${\bf H}_{n}=:(H_{n,1}),\dots,H_{n,q})$ for each $n$. Let ${\bf\overline{H}}:=(\overline{H_{1}},\dots,\overline{H_{q}})$, where ${\bf H}=:(H_{1},\dots,H_{q})\in\mathcal{P}(X)^{q}$. Let $\mathcal{K}(X)$ be the set of all elements of $\mathcal{P}(X)$ which are compact.
@@
The existence and uniqueness of the invariant set would be shown by following Hutchinson’s approach [13] and extending a result by Kocsis and Páles [15, Theorem 17] to the graph-directed framework. However, the approximation of the invariant set by the iterated images of a finite set is important in some examples including the functional equation in Section 5.
The rest of this paper is organized as follows. In Section 2, we give some preliminary results needed for the proof. Section 3 is devoted to the proof of Theorem 1.1. In Section 4, we consider transforms of semi-metrics and give an example of the graph-directed invariant set. We deal with a graph-direted de Rham’s functional equations in Section 5.
@@ 2.1. Symmetric pre-metric space
In the sequel, we need to deal with a symmetric, pre-metric space. Let $X$ be a non-empty set. We say that a map $d:X^{2}\to[0,\infty)$ is a symmetric pre-metric on $X$ if $d(x,x)=0$ for every $x\in X$, and $d(x,y)=d(y,x)$ for every $x,y\in X$. A symmetric pre-metric $d:X^{2}\to[0,\infty)$ is a semi-metric on $X$ if and only if the positivity holds, specifically, the property that $d(x,y)=0$ implies $x=y$. For a symmetric pre-metric space, we introduce a topology in the same manner as in the semi-metric case. The notions for convergence and Cauchy sequence are also defined in the same manner. However, the limit of convergent sequence may not be unique. We can also define the basic triangle function in the same manner. The regular and normal properties are also defined in the same manner.
We now state some consequences of the normal property. We say that a subset $H$ of $X$ is bounded if there exists $x_{0}\in X$ and $r_{0}>0$ such that $H\subset B(x_{0},r_{0})$. Let $\textup{diam}(H):=\sup\{d(x,y)|x,y\in A\}$ for $H\subset X$, $H\neq\emptyset$, and we call it the diameter of $H$.
Lemma 2.1 ([4, Proposition 3]).
@@
(i) Let $(x_{n})_{n=0}^{\infty}$ be a Cauchy sequence. Since $d(x_{i},x_{j})\leq\Phi_{d}(d(x_{i},x_{0}),d(x_{j},x_{0}))$, the coordinate-wise increasing property of $\Phi_{d}$, and the normal property for $(X,d)$, it suffices to show that $\sup_{i}d(x_{i},x_{0})<\infty$. By the assumption, there exists $N$ such that for every $m,n\geq N$, $d(x_{m},x_{n})\leq 1$. Then, for every $i\geq N$, $d(x_{i},x_{0})\leq\Phi_{d}(d(x_{i},x_{N}),d(x_{N},x_{0}))\leq\Phi_{d}(1,d(x_{N},x_{0}))$. Hence, $\sup_{i}d(x_{i},x_{0})<\infty$.
(ii) By Lemma 2.1, there exist $x_{0}\in X$ and $r_{0}>0$ such that $H\subset B(x_{0},r_{0})$. Let $y\in H(r)$. Then, there exists $z\in H$ such that $d(y,z)<r$. Hence, by the coordinate-wise increasing property of $\Phi_{d}$, $d(x_{0},y)\leq\Phi_{d}(d(x_{0},z),d(z,y))\leq\Phi_{d}(r_{0},r)$. By the normal property for $(X,d)$, $\Phi_{d}(r_{0},r)<\infty$. Hence, $y\in B(x_{0},\Phi_{d}(r_{0},r)+1)$. Hence, $H(r)$ is bounded.
(iii) If $H_{1}$ or $H_{2}$ is empty, then, the assertion clearly holds. Assume $H_{1}$ and $H_{2}$ are both non-empty. Let $x_{1}\in H_{1}$ and $x_{2}\in H_{2}$. By the assumption, there exist $r_{1},r_{2}>0$ such that $H_{i}\subset B(x_{i},r_{i})$, $i=1,2$. Let $y\in H_{2}$. Then, $d(x_{1},y)\leq\Phi_{d}(d(x_{1},x_{2}),d(x_{2},y))\leq\Phi_{d}(d(x_{1},x_{2}),r_{2})<\infty$. Hence, $y\in B(x_{1},\Phi_{d}(d(x_{1},x_{2}),r_{2})+1)$. Hence, $H_{1}\cup H_{2}\subset B(x_{1},r_{1}+\Phi_{d}(d(x_{1},x_{2}),r_{2})+1)$. This means that $H_{1}\cup H_{2}$ is bounded. ∎
@@ 2.2. Contraction and fixed point
(ii) If $\phi_{1}$ and $\phi_{2}$ are right-continuous comparison functions and let $\phi:=\max\{\phi_{1},\phi_{2}\}$, then, $\phi$ is also a right-continuous comparison function.
@@
Jachymski [14] gave useful equivalent conditions for Matkowski’s weak contraction. Leśniak, Snigireva and Strobin’s survey [17] contains various results for weakly contractive IFSs.
@@
Theorem 2.4 (Bessenyei and Páles [1, Theorem 1] ).
@@
We give a proof of this assertion, because in the sequel we use the following argument to establish an assertion. We adopt the strategy of the proof of [1, Theorem 1] with minor changes.
@@
Hence, $T^{n(\epsilon)}(B(x_{N(\epsilon)},\epsilon))\subset B(x_{N(\epsilon)},\epsilon)$. Hence, for every $k\geq n(\epsilon)$,
@@
By the completeness, we have the limit $x_{\infty}\in X$ such that $\lim_{n\to\infty}d(x_{n},x_{\infty})=0$.
@@
Hence, $d(Tx_{\infty},x_{\infty})=0$. Hence, by the positivity of the semi-metric $d$, $Tx_{\infty}=x_{\infty}$. ∎
@@
2.3. Kuratowski non-compactness measure
@@
where we let $\inf\emptyset:=\infty$. We call this the Kuratowski non-compactness measure of $H$. For metric spaces, this notion was introduced by [16], and, the monograph by Granas and Dugundji [11] gives properties of the Kuratowski non-compactness measure.
By the definition, $\chi(H)<\infty$ if $H$ is bounded. It holds that
@@ 2.4. Hausdorff-Pompeiu distance
Let $(X,d)$ be a a normal, regular semi-metric space. Let $\mathcal{F}(X)$ be the set of non-empty bounded subsets of $X$. We do not require that the subsets are closed or compact.
@@
(i) By the definition, $d_{HP}$ is symmetric and non-negative, specifically, $d_{HP}(A,B)=d_{HP}(B,A)\geq 0$. Furthermore, $d_{HP}(A,A)=0$.
@@
Step 2. Assume that $A,B,C\in\mathcal{F}(X)$, $r_{1}>d_{HP}(A,B)$ and $r_{2}>d_{HP}(B,C)$. Then, by Step 1, $d_{HP}(A,C)\leq\Phi_{d}(r_{1},r_{2})+\epsilon$ for every $\epsilon>0$, and hence, $d_{HP}(A,C)\leq\Phi_{d}(r_{1},r_{2})$. Hence, for every $u,v\geq 0,\epsilon>0$, $\Phi_{d_{HP}}(u,v)\leq\Phi_{d}(u+\epsilon,v+\epsilon)$. By the normal property for $d$, $\Phi_{d_{HP}}(u,v)\leq\Phi_{d}(u+\epsilon,v+\epsilon)<\infty$. Hence $d_{HP}$ satisfies the normal property.
@@
Even if $A$ is closed, we do not assure that $f(A)$ is closed. We need additional conditions for $f$. Hence, we have not assumed that any element of $\mathcal{F}(X)$ is a closed subset of $X$. To our knowledge, it is not clear whether $T(A)$ and $T(B)$ are closed in the proof of [3, Lemma 5]. The argument holds if we replace $T(A)$ and $T(B)$ with $\overline{T(A)}$ and $\overline{T(B)}$, respectively.
@@
Assume that $d_{HP}(A,B)=0$. Then, for every $n\geq 1$, $A\subset B(1/n)$. Let $x\in A$. Then, for each $n\geq 1$, there exists $y_{n}\in B$ such that $d(x,y_{n})<1/n$. By [7, Theorem 4.2], every convergent sequence in a closed subset, its limit is also in the closed subset [^1] In [4, Section 2], it is stated that the neighborhood topology and the sequential topology may differ from each other, but this is not compatible with [7, Theorem 4.2].. Since $B$ is closed, $x\in B$. Hence, $A\subset B$. In the same manner, we see that $B\subset A$. ∎

### v2

@@ Abstract.
We present a construction of graph-directed invariant sets of weak contractions in the sense of Matkowski-Rus on semi-metric spaces. We follow the approach by Bessenyei and Pénzes, which applies the Kuratowski noncompactness measure without relying on Blascke’s completeness theorem. We also establish a relationship between this approach and a generalized de Rham’s functional equation indexed by a finite directed graph.
@@ 1. Introduction
We now assume that a finite directed multigraph allowing self-loops and a metric space are given, with a transformation on the metric space assigned to each directed edge. This setting allows us to define a natural analogue of Hutchinson’s self-similar sets, known as graph-directed fractal sets. These are described as vectors of non-empty compact subsets and can be realized as fixed points of certain transformations on product spaces. See (1.1) below for the precise definition of the transformation. The fixed point is called a graph-directed invariant set, which is a notion first introduced by Mauldin and Williams [21] as a generalization of Hutchinson’s self-similar sets. When the underlying graph consists of a single vertex, the framework reduces to the classical IFS. Similar to self-similar sets, graph-directed invariant sets and related topics have been studied extensively. For further details, see the book by Mauldin and Urbański [20].
In this paper, we construct an invariant vector of non-empty compact subsets for graph-directed IFSs consisting of weak contractions on semi-metric spaces. Semi-metric is obtained by omitting the triangle inequality from Fréchet’s definition of metric. We work specifically with regular semi-metrics introduced by Bessenyei and Páles [2]. Informally, the local properties of semi-metric spaces are similar to those of metric spaces, on the other hand, a semi-metric can differ significantly in its global structure. We assign weak contractions in the sense of Matkowski-Rus [19, 25] as the components of the graph-directed IFS. Since weak contractions are less restrictive than standard contractions, a more careful analysis is required. Our arguments depend on the fixed point theorem provided in [2, Theorem 1].
Our approach follows the method introduced by Bessenyei and Pénzes [3, 4, 5], which differs from Hutchinson’s approach. The method of Bessenyei and Pénzes avoids the use of Banach’s fixed point theorem which depends on Blaschke’s completeness theorem [6], and instead applies Kuratowski’s measure of noncompactness [17]. This enables us to approximate the unique invariant non-empty compact subset by an explicitly constructible, increasing sequence of finite sets. Furthermore, it clearly demonstrates that the invariant subset depends on the topology, rather than the metric.
@@ 1.1. Framework and main result
Let $X$ be a non-empty set. We say that a map $d:X^{2}\to[0,\infty)$ is a semi-metric on $X$ if $d(x,y)=0$ if and only if $x=y$, and $d(x,y)=d(y,x)$ for every $x,y\in X$. Let $B(x,r):=\{y\in X|d(x,y)<r\}$ for $x\in X$ and $r>0$. We say that a subset $U$ of $X$ is an open set if for every $x\in X$, there exists $r>0$ such that $B(x,r)\subset U$. This defines a topology on $X$. We remark that $B(x,r)$ can be a non-open set. A closed set is a complement of an open set. For $H\subset X$, we denote the closure of $H$ by $\overline{H}$. We say that a sequence $(x_{n})_{n}$ in $X$ is convergent if $\lim_{n\to\infty}d(x_{n},x)=0$ for some $x\in X$, and it is a Cauchy sequence if $\lim_{n,m\to\infty}d(x_{n},x_{m})=0$. We say that $(X,d)$ is complete if every Cauchy sequence is convergent.
We introduce a notion which plays the role of the classical triangle inequality. For $u,v\in[0,\infty)$, let
@@
We say that a semi-metric space $(X,d)$ satisfies the regular property for the basic triangle function if $\Phi_{d}$ is continuous at $(0,0)$, and, it satisfies the normal property for the basic triangle function if $\Phi_{d}(u,v)\neq\infty$ for $u,v\in[0,\infty)$. Every metric space is a semi-metric space satisfying the regular and normal properties for the basic triangle function. The regular property was introduced in [2] and it plays the role of the classical triangle inequality. Informally, we can say that a regular metric satisfies a kind of triangle inequality in “small” scales. The normal property was introduced in [5] and it is important when we deal with bounded subsets. The regular or normal property is different from the notion of the regular or normal topological space (${\bf T}_{3}$ space or ${\bf T}_{4}$ space).
@@
We adopt Matkowski-Rus’s definition [^1] This is sometimes called the Matkowski contraction, however, the fixed point theorem for this weak contraction was independently proved by Matkowski and Rus in 1975. of weak contractions [19, 25]. We say that a map $\phi:[0,\infty)\to[0,\infty)$ is a comparison function if $\phi$ is increasing, and $\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$. Let $\phi$ be a comparison function. We say that a map $T:X\to X$ is a $\phi$-contraction if $d(Tx,Ty)\leq\phi(d(x,y))$ for every $x,y\in X$.
We state the framework of graph-directed IFSs as in [11, Section 3.1]. Let $q\geq 1$ and $V:=\{1,\cdots,q\}$. For $i,j\in V$, let $E_{i,j}$ be a finite set of directed edges from $i$ to $j$. We remark that $i=j$ is allowed. We allow multiple edges and self-loops in order to make the setting here a generalization of the self-similar set, which corresponds to the case that $q=1$. We assume that for every $i\in V$, there exists $j\in V$ such that $E_{i,j}\neq\emptyset$. This assumption is weaker than the transitivity condition. Assume that for each edge $e\in E_{i,j}$, a right-continuous comparison function $\phi_{e}$ and a $\phi_{e}$-contraction $f_{e}:X\to X$ are given. The case that $q=1$ gives the framework of Hutchinson’s self-similar sets.
Let $\mathcal{P}(X)$ be the set of non-empty subsets of $X$. Hereafter, we need to deal with product spaces $X^{q}$ and $\mathcal{P}(X)^{q}$. Elements in these product spaces are written in bold face.
@@
For a family $({\bf H}_{n})_{n}$ in $\mathcal{P}(X)^{q}$, let $\cup_{n}{\bf H}_{n}:=\left(\cup_{n}H_{n,1},\dots,\cup_{n}H_{n,q}\right)$, where ${\bf H}_{n}=:\left(H_{n,1},\dots,H_{n,q}\right)$ for each $n$. Let ${\bf\overline{H}}:=\left(\overline{H_{1}},\dots,\overline{H_{q}}\right)$, where ${\bf H}=:(H_{1},\dots,H_{q})\in\mathcal{P}(X)^{q}$. Let $\mathcal{K}(X)$ be the set of all elements of $\mathcal{P}(X)$ which are compact.
@@
The existence and uniqueness of the invariant set will be shown following Hutchinson’s approach [14] and extending a result by Kocsis and Páles [16, Theorem 17] to the graph-directed framework. However, the approximation of the invariant set by the iterated images of a finite set is important in some examples including the functional equation in Section 5.
The rest of this paper is organized as follows. In Section 2, we give some preliminary results needed for the proof. Section 3 is devoted to the proof of Theorem 1.1. In Section 4, we consider transforms of semi-metrics and give an example of the graph-directed invariant set. We deal with graph-directed de Rham’s functional equations in Section 5.
@@ 2.1. Symmetric pre-metric space
In the sequel, we need to deal with a symmetric, pre-metric space. Let $X$ be a non-empty set. We say that a map $d:X^{2}\to[0,\infty)$ is a symmetric pre-metric on $X$ if $d(x,x)=0$ for every $x\in X$, and $d(x,y)=d(y,x)$ for every $x,y\in X$. A symmetric pre-metric $d:X^{2}\to[0,\infty)$ is a semi-metric on $X$ if and only if the positivity holds, specifically, the property that $d(x,y)=0$ implies $x=y$. For a symmetric pre-metric space, we introduce a topology in the same manner as in the semi-metric case. The notions of convergence and Cauchy sequence are also defined in the same manner. However, the limit of a convergent sequence may not be unique. We can also define the basic triangle function in the same manner. The regular and normal properties are also defined in the same manner.
We now state some consequences of the normal property. We say that a subset $H$ of $X$ is bounded if there exist $x_{0}\in X$ and $r_{0}>0$ such that $H\subset B(x_{0},r_{0})$. Let $\textup{diam}(H):=\sup\{d(x,y)|x,y\in A\}$ for $H\subset X$, $H\neq\emptyset$, and we call it the diameter of $H$.
Lemma 2.1 ([5, Proposition 3]).
@@
(i) Let $(x_{n})_{n=0}^{\infty}$ be a Cauchy sequence. Since $d(x_{i},x_{j})\leq\Phi_{d}(d(x_{i},x_{0}),d(x_{j},x_{0}))$, for the coordinate-wise increasing property of $\Phi_{d}$, and the normal property for $(X,d)$, it suffices to show that $\sup_{i}d(x_{i},x_{0})<\infty$. By the assumption, there exists $N$ such that for every $m,n\geq N$, $d(x_{m},x_{n})\leq 1$. Then, for every $i\geq N$, $d(x_{i},x_{0})\leq\Phi_{d}(d(x_{i},x_{N}),d(x_{N},x_{0}))\leq\Phi_{d}(1,d(x_{N},x_{0}))$. Hence, $\sup_{i}d(x_{i},x_{0})<\infty$.
(ii) By Lemma 2.1, there exist $x_{0}\in X$ and $r_{0}>0$ such that $H\subset B(x_{0},r_{0})$. Let $y\in H(r)$. Then, there exists $z\in H$ such that $d(y,z)<r$. Hence, by the coordinate-wise increasing property of $\Phi_{d}$, $d(x_{0},y)\leq\Phi_{d}(d(x_{0},z),d(z,y))\leq\Phi_{d}(r_{0},r)$. By the normal property for $(X,d)$, $\Phi_{d}(r_{0},r)<\infty$. Hence, $y\in B(x_{0},\Phi_{d}(r_{0},r)+1)$. Thus, $H(r)$ is bounded.
(iii) If $H_{1}$ or $H_{2}$ is empty, then, the assertion clearly holds. Assume $H_{1}$ and $H_{2}$ are both non-empty. Let $x_{1}\in H_{1}$ and $x_{2}\in H_{2}$. By the assumption, there exist $r_{1},r_{2}>0$ such that $H_{i}\subset B(x_{i},r_{i})$, $i=1,2$. Let $y\in H_{2}$. Then, $d(x_{1},y)\leq\Phi_{d}(d(x_{1},x_{2}),d(x_{2},y))\leq\Phi_{d}(d(x_{1},x_{2}),r_{2})<\infty$. Hence, $y\in B(x_{1},\Phi_{d}(d(x_{1},x_{2}),r_{2})+1)$. Thus, $H_{1}\cup H_{2}\subset B(x_{1},r_{1}+\Phi_{d}(d(x_{1},x_{2}),r_{2})+1)$. This means that $H_{1}\cup H_{2}$ is bounded. ∎
@@ 2.2. Contraction and fixed point
(ii) If $\phi_{1}$ and $\phi_{2}$ are right-continuous comparison functions and $\phi:=\max\{\phi_{1},\phi_{2}\}$, then, $\phi$ is also a right-continuous comparison function.
@@
Jachymski [15] gave useful equivalent conditions for Matkowski-Rus’s weak contraction. See Berlinde, Petrusel and Rus [1] for terminology in fixed point theory. Leśniak, Snigireva and Strobin’s survey [18] contains various results for weakly contractive IFSs.
@@
Theorem 2.4 (Bessenyei and Páles [2, Theorem 1] ).
@@
We give a proof of this assertion, because in the sequel we use the following argument to establish an assertion. We adopt the strategy of the proof of [2, Theorem 1] with minor changes.
@@
Hence, $T^{n(\epsilon)}(B(x_{N(\epsilon)},\epsilon))\subset B(x_{N(\epsilon)},\epsilon)$. Thus, for every $k\geq n(\epsilon)$,
@@
By the completeness, we have the limit $x_{\infty}\in X$ such that $\lim_{n\to\infty}d(x_{n},x_{\infty})=0$. Therefore,
@@
Hence, $d(Tx_{\infty},x_{\infty})=0$. So, by the positivity of the semi-metric $d$, $Tx_{\infty}=x_{\infty}$. ∎
@@
2.3. Kuratowski measure of non-compactness
@@
where we let $\inf\emptyset:=\infty$. We call this the Kuratowski measure of non-compactness of $H$. For metric spaces, this notion was introduced in [17], and, the monograph by Granas and Dugundji [12] gives properties of the Kuratowski measure of non-compactness.
By definition, $\chi(H)<\infty$ if $H$ is bounded. It holds that
@@ 2.4. Hausdorff-Pompeiu distance
Let $(X,d)$ be a normal, regular semi-metric space. Let $\mathcal{F}(X)$ be the set of non-empty bounded subsets of $X$. We do not require that the subsets are closed or compact.
@@
(i) By definition, $d_{HP}$ is symmetric and non-negative, specifically, $d_{HP}(A,B)=d_{HP}(B,A)\geq 0$. Furthermore, $d_{HP}(A,A)=0$.
@@
Step 2. Assume that $A,B,C\in\mathcal{F}(X)$, $r_{1}>d_{HP}(A,B)$ and $r_{2}>d_{HP}(B,C)$. Then, by Step 1, $d_{HP}(A,C)\leq\Phi_{d}(r_{1},r_{2})+\epsilon$ for every $\epsilon>0$, and hence, $d_{HP}(A,C)\leq\Phi_{d}(r_{1},r_{2})$. Thus, for every $u,v\geq 0,\epsilon>0$, $\Phi_{d_{HP}}(u,v)\leq\Phi_{d}(u+\epsilon,v+\epsilon)$. By the normal property for $d$, $\Phi_{d_{HP}}(u,v)\leq\Phi_{d}(u+\epsilon,v+\epsilon)<\infty$. Hence $d_{HP}$ satisfies the normal property.
@@
Even if $A$ is closed, we do not assure that $f(A)$ is closed. Let $X=(0,\infty)\setminus\{1\}$ be equipped with the topology induced by the Euclidean metric. Let $A:=(0,1)\subset X$. Then, it is a closed subset of $X$. Let $f(x)=(x/2)+1$. Then, $f(X)=(1,\infty)\setminus\{3/2\}\subset X$, and $f(A)=(1,3/2)$, which is not closed in $X$. Therefore, $T(A)$ and $T(B)$ may not be closed in the proof in [4, Lemma 5]. The argument holds if we replace $T(A)$ and $T(B)$ with $\overline{T(A)}$ and $\overline{T(B)}$, respectively. Therefore, we have not assumed that any element of $\mathcal{F}(X)$ is a closed subset of $X$.
@@
Assume that $d_{HP}(A,B)=0$. Then, for every $n\geq 1$, $A\subset B(1/n)$. Let $x\in A$. Then, for each $n\geq 1$, there exists $y_{n}\in B$ such that $d(x,y_{n})<1/n$. By [8, Theorem 4.2], the limit of every convergent sequence in a closed subset is also in the closed subset [^2] In [5, Section 2], it is stated that the neighborhood topology and the sequential topology may differ from each other, but this is not compatible with [8, Theorem 4.2].. Since $B$ is closed, $x\in B$. Hence, $A\subset B$. In the same manner, we see that $B\subset A$. ∎